# Exploring Training and Inference Scaling Laws in Generative Retrieval

### v1

@@ 3.3.2. Training Setting
$\bullet$ N-Gram-Based Generative Retrieval. For n-gram-based generative retrieval, we use the Natural Questions (NQ) (Kwiatkowski et al., 2019) dataset, which contains over 20 million documents. Following the MINDER methodology, which originally involves three types of identifiers, we focus exclusively on the body text as the identifier type. This simplification is made because our study emphasizes scaling behavior rather than achieving absolute performance, and the body text serves as the most important identifier type. For each document, we select 10 n-grams, each consisting of 10 tokens, based on their overlap with the associated query to ensure high semantic relevance. The final training set consists of nearly 600,000 query-to-n-gram pairs.
@@
where $n$ represents an individual n-gram and $\theta$ represents the model parameters. Since each n-gram is treated independently, the model learns to predict each query-relevant n-gram as a separate target during training.
To keep consistent with the following setup of codebook-based methods, we train each model for one epoch using the MINDER-provided data with LoRA (Hu et al., 2022). As recommended in MINDER, the learning rate for T5 models is set to 3e-5. For LLaMA models, we use a learning rate of 3e-4, a commonly adopted value for fine-tuning LLaMA.
$\bullet$ Codebook-Based Generative Retrieval. Following the selected representative codebook-based generative retrieval method, RIPOR (Zeng et al., 2024), we conduct experiments using the MSMARCO-1M dataset, a subset of the MSMARCO (Nguyen et al., 2017) dataset, containing one million passages and query-document pairs. This dataset is chosen because RIPOR is only available on MSMARCO, and since our study focuses on scaling behavior rather than absolute performance, the choice of the dataset has minimal impact on the relative results.
@@ 3.4. Evaluation
In generative retrieval, the primary objective is to generate identifiers that correspond to relevant documents while distinguishing them from irrelevant ones. To quantify this ability, we define a contrastive generation loss (CGL), which evaluates the model’s capacity to generate identifiers for positive documents in the presence of negative documents.
@@
The contrastive generation loss is a continuous and highly sensitive metric designed to effectively evaluate retrieval performance across different generative retrieval methods and models. 1) For n-gram-based generative retrieval, where each document is represented by multiple identifiers, the $\mathcal{L}(q,\mathcal{I}_{d^{+}})$ is extended by averaging the generation loss across all n-grams associated with the positive document. 2) For codebook-based generative retrieval, the query-to-document relationship is evaluated through the loss of generating the document’s unique code sequence. 3) Furthermore, this metric is versatile across different LLM architectures, as it directly leverages the generative capabilities of the models to compute losses. 4) Moreover, the CGL is based on the relative proportion of the positive document’s loss compared to negative documents, reducing sensitivity to the absolute magnitude of the loss and ensuring consistency across varying model scales and training settings.
To validate the proposed CGL, we conducted experiments to analyze its relationship with existing ranking metrics, such as Recall, NDCG, MAP, and MRR. Using the MINDER framework as an example, we trained models of varying sizes and configurations and then calculated their retrieval performance using Recall@100, NDCG@10, MAP@10, MRR@10 as well as the proposed CGL. Results in Figure 2 reveal an almost linear relationship between contrastive generation loss and standard metrics. This alignment demonstrates that the proposed metric effectively captures retrieval performance, offering a consistent and reliable evaluation framework across different settings.
@@ 4. Training Scaling Laws
In this section, we present the results of our experiments and summarize our initial investigation into the scaling laws for generative retrieval. Specifically, we analyze how model size, training data size, and identifier methods influence retrieval performance, using contrastive generation loss as the evaluation metric. These experiments aim to uncover the relationships and trends that define the scalability of generative retrieval systems.
@@ 4.1. Model Size Scaling
We now investigate the impact of model size on generative retrieval performance. We begin by examining n-gram-based generative retrieval.
@@ 4.1.1. N-Gram-Based Generative Retrieval
We investigated the effect of model size on generative retrieval performance using the n-gram-based method. Models of varying sizes, including T5 and LLaMA, were fine-tuned on query-to-n-gram training pairs, and their performance was evaluated using the contrastive generation loss on the test set. Figure 3 (1) illustrates the scaling behavior of T5 and LLaMA models concerning this metric. Based on the observed relationship between model size and the contrastive generation loss, we propose a scaling law to quantify this behavior as follows:
Here, $P$ represents the number of non-embedding parameters of the model, and $\mathcal{L}_{\text{CGL}}(P)$ denotes the contrastive generation loss on the test set. The parameters $\gamma$, $\alpha$, and $\lambda_{P}$ are coefficients determined through fitting. Here, $\lambda_{P}$ represents the irreducible loss, a theoretical lower bound on performance as $P$ approaches infinity, accounting for limitations such as dataset noise and variability in relevance judgments.
Using the least squares method, we fit the scaling law and report the fitted coefficients for T5 and LLaMA in Table 1. The results reveal several important insights: 1) Both models demonstrate a strong power-law relationship between model size and contrastive generation loss, with exceptionally high coefficients of determination (T5: R² = 0.996, LLaMA: R² = 0.999). 2) LLaMA demonstrates comprehensive performance advantages over T5, characterized by a more efficient scaling mechanism. Specifically, LLaMA achieves lower contrastive generation loss across model sizes and exhibits a steeper improvement curve (scaling exponent of $\alpha=2.40$ versus T5’s $\alpha=0.40$). Notably, LLaMA’s smaller irreducible loss ($\lambda_{P}$) suggests a higher potential performance ceiling, indicating its superior capability to approach the theoretical limits of generative retrieval performance. These findings highlight LLaMA’s promising performance and potentially signal the broader potential of decoder-only architectures.
@@ 4.1.2. Codebook-Based Generative Retrieval
Despite the lack of scaling trends, LLaMA consistently achieves lower contrastive generation loss than T5 across all model sizes, highlighting its stronger retrieval capabilities. These findings suggest LLaMA demonstrates promising performance characteristics and may signal avenues for future research in generative retrieval.
@@
The results presented in Figure 3 highlight clear differences between the n-gram-based and codebook-based methods in terms of scaling behavior and overall performance. 1) The n-gram-based method significantly outperforms the codebook-based approach, demonstrating substantially lower contrastive generation loss. Even the most advanced LLaMA models utilizing codebook tokens cannot match the performance of T5 models with n-gram-based retrieval. This performance gap highlights the intrinsic challenges of codebook tokens, which lack the semantic coherence and natural language alignment inherent in n-grams. 2) LLaMA consistently outperforms T5 across both methods, achieving lower CGL at comparable model sizes. This highlights LLaMA’s stronger generative capabilities and its architectural advantage.
@@ 4.2. Data Size Scaling
4.2.1. N-Gram-Based Generative Retrieval
Here, $D$ represents the number of query-identifier pairs, and $\mathcal{L}_{\text{CGL}}(D)$ denotes the contrastive generation loss on the test set. The parameters $\eta$, $\beta$, and $\lambda_{D}$ are coefficients determined through fitting. The term $\lambda_{D}$ represents the irreducible loss, a theoretical lower bound on retrieval performance as $D$ approaches infinity.
Using the least squares method, we fit the scaling law to the observed data, achieving a coefficient of determination of $R^{2}=0.990$, which indicates a strong fit. As seen in Figure 4 (1), retrieval performance improves significantly as the training data size increases, with the contrastive generation loss decreasing sharply. The power-law scaling behavior reflects the model’s capacity to leverage larger datasets to better capture the semantic relationships between queries and identifiers.
@@ 4.2.2. Codebook-Based Generative Retrieval
Using the same power-law equation as for n-grams, we fit the scaling behavior of the codebook-based method. The fitted curve in Figure 4 (2) achieves a coefficient of determination of $R^{2}=0.983$, indicating a strong fit. As the training data size increases, the contrastive generation loss decreases steadily, demonstrating that retrieval performance improves with larger datasets. The results highlight that even for the codebook-based method, which involves learning entirely new representations unrelated to the model’s pretraining objectives, increasing the data size leads to significant retrieval performance enhancements.
@@
For n-gram-based methods, the scaling exponent ($\beta=3.99$) is much larger than that of codebook-based methods ($\beta=1.01$), indicating a steeper improvement in performance with increased data size. This can be attributed to the semantic richness of n-grams, which align closely with the model’s pretraining objectives, allowing the model to fully leverage larger datasets. In contrast, the codebook-based method lacks such alignment, resulting in a slower rate of improvement as data size increases. The low scaling exponent implies that this method requires substantially more training data to achieve comparable performance. Recent studies suggest more advanced training strategies, such as ranking losses (Tang et al., 2024; Zeng et al., 2024), could potentially address these learning challenges.
@@ 5. Inference Scaling Laws
In the previous sections, we demonstrated the existence of scaling laws in generative retrieval concerning model size and data size, highlighting how larger models and richer data yield stronger performance. These findings align with trends observed in LLMs, further validating the synergy between generative retrieval and LLMs.
Recent studies have also revealed another dimension of scaling: the scaling of computational investment during inference (Wu et al., 2024). Specifically, increasing inference computing, such as scaling the number of decoding tokens, has been shown to yield substantial performance gains. This trend suggests that inference scaling may be a critical factor in optimizing LLM-based systems, complementing improvements from larger models and data.
In generative retrieval, inference scaling is particularly promising because the identifiers used for retrieval are generated dynamically at this stage. Generative retrieval inherently involves generating query-specific identifiers to score and rank documents. The quality and diversity of generated identifiers are directly influenced by the inference process, including parameters such as beam size, which controls the breadth of the search during generation. Larger beam sizes allow for the exploration of a broader search space, potentially improving retrieval performance, but they also increase inference latency, creating a trade-off between quality and efficiency.
Motivated by this, we investigate whether the inference scaling laws observed in LLMs extend to generative retrieval. Given that generative retrieval shares a strong alignment with the LLMs, we wonder whether similar scaling behaviors will hold for generative retrieval inference.
@@ 5.1. Experimental Setup
We use the n-gram-based generative retrieval method described in Section 3.2.1. And we chose T5-Base and LLaMA-7B as representative models, both are fine-tuned under identical settings on the NQ dataset (Kwiatkowski et al., 2019), using the same query-to-n-gram pairs. Each model is trained for one epoch to ensure consistency across experiments.
During inference, we vary the beam size, testing values of $B=\{1,5,10,20,50,100\}$ and record the corresponding inference computation needed. Beam size determines the number and search space of candidate n-grams generated per query, which in turn affects the diversity of identifiers used for document retrieval. Increasing the beam size effectively increases the computational cost during inference, as a larger search space requires more floating-point operations (FLOPs) to generate candidate n-grams. To ensure comparability, the number of retrieved documents is fixed at 100 for all configurations, and other settings like batch size remain constant.
To evaluate retrieval performance, we define a metric called Miss Rate (MR), which measures the proportion of relevant documents that are not retrieved within the top $k$ results.
@@ 5.2. Results
The results of our experiments are summarized in Figure 5, which illustrates the Miss Rate across different inference FLOPs per query for both T5-Base and LLaMA-7B models. We evaluate MR at different retrieval thresholds ($k=5,20,100$) to assess the retrieval performance under varying levels of precision.
@@
The fitted curves in Figure 5, reveal a consistent trend for both T5-Base and LLaMA-7B models: the Miss Rate decreases steadily as the inference FLOPs per query increase. This validates the effectiveness of inference scaling in improving generative retrieval performance. The decrease in MR is particularly pronounced at lower computational budgets, where increased beam sizes lead to significant performance gains.
As shown in Table 3, the proposed scaling law provides an excellent fit to the experimental data for both models, with high $R^{2}$ values in most cases. Additionally, the rate of decline in MR varies with the retrieval threshold $k$. For both models, MR@5 decreases the fastest, reflecting the substantial benefits of inference scaling for high-precision retrieval. Conversely, MR@100 exhibits a slower decline, indicating that broader recall settings are less sensitive to increased computational resources.
Comparing the two models, when inference FLOPs are below $10^{11}$, T5-Base outperforms LLaMA-7B, achieving lower Miss Rates across all $k$ values. However, as FLOPs increase beyond $10^{11}$, LLaMA-7B demonstrates a clear advantage, with all three Miss Rates achieving lower values. This highlights that LLaMA-7B gains substantially more performance rise from increased inference computation compared to T5-Base, making it highly effective in high-computation scenarios. Notably, LLaMA-7B’s irreducible loss ($\lambda_{C}$) for MR@5 is significantly lower than T5-Base, indicating superior potential for high-precision retrieval, while its irreducible loss for MR@20 and MR@100 is slightly higher, suggesting T5-Base retains an advantage in broader recall settings. These results highlight the trade-offs between the two models, with LLaMA-7B excelling in high-precision tasks, achieving lower Miss Rates for the top-ranked results as computational budgets grow, while T5-Base demonstrates greater potential for broader recall tasks like larger-scale retrieval.
@@ 6. Discussion
In this work, we explored the training scaling laws of model size, data size, and inference scaling laws in generative retrieval, proposing metrics to analyze retrieval performance under varying computational budgets. By systematically varying these factors, we highlighted how scaling influences retrieval performance and identified key trade-offs between inference computation and performance. While our findings provide valuable insights into generative retrieval systems, several important aspects remain unexplored and warrant further investigation.
@@
We introduced a new metric contrastive generation loss to evaluate the generative retrieval capability of models. While CGL provides a relative measure of a model’s retrieval ability under the same settings, it is not equivalent to traditional metrics such as Recall. For example, two models with similar CGL values may exhibit differing recall performance, and a lower CGL does not guarantee higher recall. CGL is useful for comparing models in relative terms within the same experimental setup, but its interpretability as an absolute measure of retrieval performance is limited.
Another limitation of this work lies in the training methodology, as we relied solely on the basic generative cross-entropy loss. While this aligns well with n-gram-based methods, which match the model’s pretraining objective and require minimal fine-tuning, it poses challenges for codebook-based methods that need to learn entirely new tokens or relationships, demanding more training data and epochs. We attempted to use RIPOR’s full dataset, containing over 80 million query-to-code sequences from MSMARCO, but this scale of data was computationally infeasible for LLaMA, as it would require thousands of hours for us to complete the training. Future research with greater resources could further explore the potential of such extensive datasets and more advanced training objectives.
@@
In our experiments, we observed that codebook-based methods did not exhibit a clear scaling effect with increasing model size, but showed a pronounced scaling effect with training data size. This discrepancy may stem from the significantly higher learning difficulty of codebook representations compared to n-grams. The amount of training data used in our experiments may have been insufficient, leaving the models in an underfitted state and preventing the scaling effect with model size from emerging. Previous studies have shown that model performance can undergo substantial improvement at a certain point (Du et al., 2024), it is possible that with sufficient training data and time, codebook-based methods might also exhibit scaling with model size. Future research with enhanced settings and resources could further investigate this possibility.
@@
In this study, we primarily focused on the n-gram-based method for inference scaling. We observed improved retrieval performance by increasing beam size and validated its alignment with power-law scaling. However, codebook-based methods could also potentially benefit from inference scaling strategies, such as increasing beam size to explore a larger search space. Due to space limitations, we did not explore this direction, but it represents a possible direction for future research.

### v2

@@ 3.3.2. Training Setting
$\bullet$ N-Gram-Based Generative Retrieval. For n-gram-based generative retrieval, we use the Natural Questions (NQ) (Kwiatkowski et al., 2019) dataset, which contains over 20 million documents. Following the MINDER methodology, which originally involves three types of identifiers, we focus exclusively on the body text as the identifier type. This simplification is made because our study emphasizes scaling behavior rather than achieving absolute performance, and the body text serves as the most important identifier type. For each document, we select 10 n-grams, each consisting of 10 tokens, based on their overlap with the associated query to ensure semantic relevance. The final training set includes nearly 600,000 query-to-n-gram pairs.
@@
where $n$ represents an individual n-gram and $\theta$ represents the model parameters. Since each n-gram is treated independently, the model learns to predict each query-relevant n-gram as a separate target.
To keep consistent with the following setup of codebook-based methods, we train each model for one epoch using the MINDER-provided data with LoRA (Hu et al., 2022). As recommended in MINDER, the learning rate for T5 models is set to 3e-5. For LLaMA models, we use a learning rate of 3e-4, a commonly adopted value.
$\bullet$ Codebook-Based Generative Retrieval. Following the selected representative codebook-based generative retrieval method, RIPOR (Zeng et al., 2024b), we conduct experiments using the MSMARCO-1M dataset, a subset of the MSMARCO (Nguyen et al., 2017) dataset, containing one million passages and query-document pairs. This dataset is chosen because RIPOR is only available on MSMARCO, and since our study focuses on scaling behavior rather than absolute performance, the choice of the dataset has minimal impact on the relative results.
@@ 3.4. Evaluation
$\bullet$ Contrastive generation loss. In generative retrieval, the primary objective is to generate identifiers that correspond to relevant documents while distinguishing them from irrelevant ones. To quantify this ability, we define a contrastive generation loss (CGL), which evaluates the model’s capacity to generate identifiers for positive documents in the presence of negative documents.
@@
Unlike contrastive loss (Khosla et al., 2020) or contrastive entropy (Fang et al., 2024) which operates in embedding space, CGL directly uses generation loss to assess the model’s preference for positive identifiers. It requires no additional training objective and works purely in the generation setting. CGL has the following advantages:
  • Method compatibility. CGL supports both n-gram-based and codebook-based generative retrieval. For n-gram identifiers, it averages the generation loss overall n-grams of the positive document. For codebook methods, it evaluates the loss of generating the complete token sequence assigned to each document.
  • Model compatibility. CGL is applicable across different LLMs, as it operates solely on forward generation loss without requiring model-specific structures or additional training objectives.
  • Relative evaluation. By using the ratio between positive and negative losses, CGL mitigates sensitivity to absolute loss values, ensuring consistent evaluation.
$\bullet$ Validation. To validate the proposed CGL, we conducted experiments to analyze its relationship with existing ranking metrics, such as Recall, NDCG, MAP, and MRR. Using the MINDER framework as an example, we trained models of varying sizes and configurations and then calculated their retrieval performance using Recall@100, NDCG@10, MAP@10, MRR@10 as well as the proposed CGL. Results in Figure 3 reveal an almost linear relationship between contrastive generation loss and standard metrics. This alignment demonstrates that the proposed metric effectively captures retrieval performance, offering a consistent and reliable evaluation framework across different settings.
@@ 4. Training Scaling Laws
In this section, we present the results of our experiments and summarize our investigation into the training scaling laws for generative retrieval. Specifically, we analyze how model size, training data size, and identifier methods influence retrieval performance, using CGL as the evaluation metric.
@@ 4.1. Model Size Scaling
We now investigate the impact of model size on retrieval performance. We begin by examining n-gram-based generative retrieval.
@@ 4.1.1. N-Gram-Based Generative Retrieval
We investigated the effect of model size on generative retrieval performance using the n-gram-based method. Models of varying sizes, including T5 and LLaMA, were fine-tuned on query-to-n-gram training pairs, and their performance was evaluated using the CGL on the test set. Figure 4 (1) illustrates the scaling behavior of T5 and LLaMA models concerning this metric. Based on the observed relationship between model size and the CGL, we propose a scaling law to quantify this behavior as follows:
Here, $P$ represents the number of non-embedding parameters of the model, and $\mathcal{L}_{\text{CGL}}(P)$ denotes the CGL on the test set. The parameters $\gamma$, $\alpha$, and $\lambda_{P}$ are coefficients determined through fitting. Here, $\lambda_{P}$ represents the irreducible loss, a theoretical lower bound on performance as $P$ approaches infinity, accounting for limitations such as dataset noise and variability in relevance judgments.
Using the least squares method, we fit the scaling law and report the fitted coefficients for T5 and LLaMA in Table 1. The results reveal several important insights: 1) Both models demonstrate a strong power-law relationship between model size and contrastive generation loss, with exceptionally high coefficients of determination. 2) LLaMA demonstrates comprehensive performance advantages over T5, characterized by a more efficient scaling mechanism. Specifically, LLaMA achieves lower CGL across model sizes and exhibits a steeper improvement curve (scaling exponent of $\alpha=2.40$ versus T5’s $\alpha=0.40$). LLaMA’s smaller irreducible loss ($\lambda_{P}$) suggests a higher potential performance ceiling, indicating its superior capability to approach the theoretical limits of generative retrieval performance. These findings highlight LLaMA’s promising performance and signal the potential of decoder-only architectures.
@@ 4.1.2. Codebook-Based Generative Retrieval
Despite the lack of scaling trends, LLaMA consistently achieves lower CGL than T5 across all model sizes, highlighting its stronger retrieval capabilities. These findings suggest LLaMA demonstrates promising performance characteristics and may signal avenues for future research in generative retrieval.
@@
The results presented in Figure 4 highlight clear differences between the n-gram-based and codebook-based methods in terms of scaling behavior and overall performance. 1) The n-gram-based method significantly outperforms the codebook-based approach, demonstrating lower CGL. Even the most advanced LLaMA models utilizing codebook tokens cannot match the performance of T5 models with n-gram-based retrieval. This performance gap highlights the intrinsic challenges of codebook tokens, which lack the semantic coherence and natural language alignment inherent in n-grams. 2) LLaMA consistently outperforms T5 across both methods, achieving lower CGL at comparable model sizes. This highlights LLaMA’s stronger generative capabilities and its architectural advantage.
@@ 4.2. Data Size Scaling
4.2.1. N-Gram-Based Generative Retrieval
Here, $D$ represents the number of query-identifier pairs, and $\mathcal{L}_{\text{CGL}}(D)$ denotes the CGL on the test set. The parameters $\eta$, $\beta$, and $\lambda_{D}$ are coefficients determined through fitting. The term $\lambda_{D}$ represents the irreducible loss, a theoretical lower bound on retrieval performance as $D$ approaches infinity.
Using the least squares method, we fit the scaling law to the observed data, achieving a coefficient of determination of $R^{2}=0.990$, which indicates a strong fit. As seen in Figure 5 (1), retrieval performance improves significantly as the training data size increases, with the CGL decreasing sharply. The power-law scaling behavior reflects the model’s capacity to leverage larger datasets to better capture the semantic relationships between queries and identifiers.
@@ 4.2.2. Codebook-Based Generative Retrieval
Using the same power-law equation as for n-grams, we fit the scaling behavior of the codebook-based method. The fitted curve in Figure 5 (2) achieves a coefficient of determination of $R^{2}=0.983$, indicating a strong fit. As the training data size increases, the CGL decreases steadily, demonstrating that retrieval performance improves with larger datasets. The results highlight that even for the codebook-based method, which involves learning entirely new representations unrelated to the model’s pretraining objectives, increasing the data size leads to performance enhancements.
@@
For n-gram-based methods, the scaling exponent ($\beta=3.99$) is much larger than that of codebook-based methods ($\beta=1.01$), indicating a steeper improvement in performance with increased data size. This can be attributed to the semantic richness of n-grams, which align closely with the model’s pretraining objectives, allowing the model to fully leverage larger datasets. In contrast, the codebook-based method lacks such alignment, resulting in a slower rate of improvement as data size increases. The low scaling exponent implies that this method requires substantially more training data to achieve comparable performance. Recent studies suggest more advanced training strategies, such as ranking losses (Tang et al., 2024; Zeng et al., 2024b), could potentially address these learning challenges.
@@ 5. Inference Scaling Laws
In the previous sections, we demonstrated the existence of training scaling laws in generative retrieval concerning model size and data size. Beyond training, recent studies have also revealed another dimension of scaling: the scaling of computational investment during inference (Wu et al., 2024). Specifically, increasing inference computing, such as scaling the number of decoding tokens, yielded substantial performance gains. This raises the question of whether similar benefits extend to generative retrieval.
Inference scaling is particularly promising in generative retrieval because the identifiers used for retrieval are generated dynamically at this stage. The generation process, controlled by parameters such as beam size, directly affects the quality and diversity of identifiers, and hence retrieval performance. Motivated by these observations, we investigate whether inference scaling laws observed in LLMs also hold in generative retrieval, and how factors like beam size and compute budget impact its effectiveness.
@@ 5.1. Experimental Setup
$\bullet$ Implementation details. We use the n-gram-based generative retrieval method described in Section 3.2.1. And we chose T5-Base and LLaMA-7B as representative models, both are fine-tuned under identical settings on the NQ dataset (Kwiatkowski et al., 2019), using the same query-to-n-gram pairs. Each model is trained for one epoch to ensure consistency across experiments.
During inference, we vary the beam size, testing values of $B=\{1,5,10,20,50,100\}$ and record the corresponding inference computation needed. Beam size determines the number and search space of candidate n-grams generated per query, affecting the diversity of identifiers used for document retrieval. Increasing the beam size effectively increases the computational cost during inference, as a larger search space requires more floating-point operations (FLOPs) to generate candidate n-grams.
$\bullet$ Evaluation. To evaluate retrieval performance, we define a metric called Miss Rate (MR), which measures the proportion of relevant documents that are not retrieved within the top $k$ results.
@@ 5.2. Results
The results of our experiments are summarized in Figure 6, which illustrates the MR across different inference FLOPs per query for both T5-Base and LLaMA-7B models. We evaluate MR at different retrieval thresholds ($k=5,20,100$) to assess the retrieval performance under varying levels of precision.
@@
The fitted curves in Figure 6, reveal a consistent trend for both T5-Base and LLaMA-7B models: the MR decreases steadily as the inference FLOPs per query increase. This validates the effectiveness of inference scaling in improving generative retrieval performance. As shown in Table 3, the proposed scaling law fits well with the experimental results. The results reveal several key findings:
  • MR@k sensitivity. The rate of MR decline varies with the retrieval threshold $k$. MR@5 decreases the fastest for both models, showing strong benefits from inference scaling in high-precision retrieval, while MR@100 declines more slowly, indicating reduced sensitivity in broader recall settings.
  • Model performance under low compute. When inference FLOPs are below $10^{11}$, T5-Base achieves lower MR across all $k$ than LLaMA-7B, suggesting that T5-Base is more efficient in compute-constrained scenarios.
  • Model performance under high compute. As FLOPs increase beyond $10^{11}$, LLaMA-7B gradually surpasses T5-Base, achieving lower MR@100 and MR@20, demonstrating greater benefits from increased inference computation.
  • Irreducible loss. LLaMA-7B exhibits lower irreducible loss ($\lambda_{C}$) for MR@5, suggesting superior capacity for high-precision retrieval. In contrast, it shows a slightly higher irreducible loss for MR@20 and MR@100, indicating that the T5-Base retains an advantage in broader recall scenarios.
@@ 6. Discussion
In this work, we explored the training and inference scaling laws in generative retrieval, proposing metrics to analyze retrieval performance under varying conditions. While our findings offer useful insights, several limitations and open questions remain.
@@
$\bullet$ Limitation of CGL. While the proposed Contrastive Generation Loss reflects the model preference for correct identifiers under consistent settings, it does not directly correlate with traditional metrics like Recall. For example, models with similar CGL values may have different recall, and a lower CGL does not guarantee better performance. CGL is effective for relative comparisons, but its interpretability as an absolute performance measure is limited.
$\bullet$ Simplified training settings. For simplicity, we adopt the standard generative cross-entropy loss without incorporating more advanced objectives. While this may be sufficient for n-gram-based methods, which match the model’s pretraining objective, it poses challenges for codebook-based methods that need to learn entirely new tokens or relationships, demanding more training data and epochs. Prior work has explored more effective alternatives, such as learning-to-rank (LTR) losses (Li et al., 2024b; Tang et al., 2024), which may better capture the retrieval objective. Future research could revisit this direction with more advanced training objectives.
@@
$\bullet$ Model scaling of the codebook-based method. In our experiments, codebook-based methods showed clear scaling with training data size but not with model size. This may be due to the greater learning difficulty of codebook representations. The training data and time in our experiments may have been insufficient, preventing the scaling effect from emerging. Previous studies have shown that model performance can undergo substantial improvement at a certain point (Du et al., 2024). Although we attempted to use RIPOR’s full dataset (over 80M query-code pairs), the computational cost for LLaMA was unaffordable. Future work could revisit this setting to better understand the scaling potential of codebook-based methods.
@@
$\bullet$ Inference scaling of the codebook-based method. Our inference scaling experiments focused on n-gram-based retrieval, showing improved performance with increased beam size and alignment with power-law scaling. Codebook methods may also benefit from inference-time strategies such as beam search (Zeng et al., 2024a), but we leave this exploration to future work due to space and resource constraints.
$\bullet$ Dataset inconsistency across methods. In our experiments, the n-gram-based and codebook-based methods were trained and evaluated on different datasets, as we directly adopted identifier sets from prior work to simplify implementation and isolate scaling effects. Consequently, the comparison between methods may be affected by dataset differences. A more controlled comparison would involve evaluating both methods on a unified dataset—a direction we leave for future work.